OpenStudy (anonymous):

what are the exact values of sin, cos, tan, sec, csc, & cot at π & 3π/2

OpenStudy (anonymous):

look at cheat sheet here. bottom page has the unit circle all mapped out http://tutorial.math.lamar.edu/cheat_table.aspx

OpenStudy (anonymous):

I was looking for something like this: http://www.analyzemath.com/trigonometry/special_angles.html

OpenStudy (anonymous):

find\[\pi\] on the unit circle. the coordinates are (-1,0) first is cosine, second is sine tangent is sine/cosine , secant is 1/cosine cosecant is 1/sine and cotangent is cosine / sin

OpenStudy (anonymous):

well that only goes from 0 to \[\frac{\pi}{2}\]

OpenStudy (anonymous):

you need the whole circle. look at the link i sent bottom page of trig cheat sheet. has the whole circle mapped

OpenStudy (anonymous):

okayyy thanks!

OpenStudy (anonymous):

once you find the coordinate you read of the answer. for example the coordinate for \[\frac{3\pi}{2}\] is (0,-1) and just read the answers from there. don't forget that 1/0 is undefined!